OpenStudy (anonymous):

simplify cosX/tan^9(X) - cosx/sin^2(x) please show method for solving

OpenStudy (mertsj):

Why did you repost this problem?

OpenStudy (mertsj):

Are you sure that exponent 9 on the tan function is correct?

OpenStudy (anonymous):

oops nope change that to a 2

OpenStudy (mertsj):

Why did you repost this problem?

OpenStudy (mertsj):

Balaa solved it before.

OpenStudy (anonymous):

oh i didnt know that i'll look

OpenStudy (mertsj):

\[\frac{cosx}{\tan ^{2}x}-\frac{cosx}{\sin ^{2}x}\]